# ACTIVA: Amortized Causal Effect Estimation without Graphs via Transformer-based Variational Autoencoder

### v1

@@ Abstract
Predicting the distribution of outcomes under hypothetical interventions is crucial in domains like healthcare, economics, and policy-making. Current methods often rely on strong assumptions, such as known causal graphs or parametric models, and lack amortization across problem instances, limiting their practicality. We propose a novel transformer-based conditional variational autoencoder architecture, named ACTIVA, that extends causal transformer encoders to predict causal effects as mixtures of Gaussians. Our method requires no causal graph and predicts interventional distributions given only observational data and a queried intervention. By amortizing over many simulated instances, it enables zero-shot generalization to novel datasets without retraining. Experiments demonstrate accurate predictions for synthetic and semi-synthetic data, showcasing the effectiveness of our graph-free, amortized causal inference approach.
Machine Learning, ICML
@@ 1 Introduction
Causal effect estimation from observational data is an important task in many domains including healthcare (Shi & Norgeot, 2022), economics (Panizza & Presbitero, 2014), and finance (Kumar et al., 2023). The task can be summarized by the following question: Given observations of certain variables, how will their distribution change when a specific action is performed?
Additionally to answering cause-effect questions, this distribution can provide further task-relevant insights such as uncertainty while preventing erroneous average effect predictions (Rissanen & Marttinen, 2021). However, distributional causal effect estimation is a challenging task, due to possible lack of interventional data, abundance of possibilities for modeling estimated distributions, and the inherent challenges of identifiability from observational data (Bareinboim et al., 2022), among others.
Figure 1: Overview of ACTIVA. Based on samples $\boldsymbol{D}^{\mathcal{M}}$ from an observational distribution $p_{\mathcal{M}}(\boldsymbol{V})$ and a query intervention $do(X=x)$, the encoder determines the latent distribution. Latent variables $\boldsymbol{z}$ are then drawn from this distribution and provided to the decoder along with the query intervention. The decoder then computes the parameters of a multivariate Gaussian representing the interventional distribution $p_{\mathcal{M}}(\boldsymbol{V}\mid do(X=x))$.
To cope with these difficulties, current approaches often rely on strong assumptions, such as having prior knowledge of the causal graph and imposing parametric assumptions on the underlying causal model (see Section 5 for more detail). For related causal tasks, there have been various proposals to tackle these challenges by extracting causal information from datasets via neural networks (Lorch et al., 2022; Scetbon et al., 2024; Sauter et al., 2024b; Annadani et al., 2025; Hollmann et al., 2025). Results from these approaches suggest that we can effectively amortize over datasets coming from different causal models for various downstream tasks. In a similar setup, amortized causal effect estimation has been studied but with restrictions to scenarios with known causal graphs (Mahajan et al., 2024).
In this paper, we overcome these restrictions by proposing a causal encoder-decoder architecture called ACTIVA. Conditioned on an interventional query, the encoder maps the observational data to a latent space and the conditional decoder transforms the encoded causal information into a mixture of Gaussians, representing the interventional distribution. Moreover, both the encoder and decoder are equivariant regarding variable ordering. We derive the evidence lower bound (ELBO) (Kingma & Welling, 2013) for our amortized learning setup with observational and interventional training data, optimizing for reconstruction of the intervened data while simultaneously regularizing the latent space.
@@
Importantly, there is no need to have access to the underlying causal graph neither during training nor inference, avoiding common pitfalls when relying on such graphs (Poinsot et al., 2024). In empirical evaluations, we show that our architecture can be trained on synthetic data and then be applied to novel instances during inference. This allows for the potential zero-shot transfer from simulated scenarios to the real-world ones. Below we provide a list of our main contributions:
@@
  • We provide a conditional encoder-decoder architecture for amortized causal effect prediction. Our model takes an observational dataset and a query intervention as input, predicts a latent representation of the dataset and outputs an estimate of the respective interventional distribution as visualized in Figure 1.
  • We formulate a loss function for amortized causal effect prediction based on the ELBO, optimizing for the conditional reconstruction of interventional data from observational data while regularizing the latent space.
@@
  • We show empirically that ACTIVA successfully predicts interventional distributions given observational data and interventional queries both on synthetic and semi-synthetic data.
@@
Overall, our work highlights the significance of amortized causal inference as a tool to overcome traditional hurdles in distributional causal effect estimation. [^1] The code to our model is available at https://anonymous.4open.science/r/Amortized_Interventional_Distribution_Estimation-30B2/.
2 Background
2.1 Background on Conditional $\beta$-VAEs
@@
Variational Autoencoders (VAEs) are generative models that approximate the joint distribution $p(\boldsymbol{x},\boldsymbol{z})=p(\boldsymbol{x}\mid\boldsymbol{z})p(\boldsymbol{z})$, where $\boldsymbol{z}$ are latent variables capturing unobserved variables governing the data. The marginal likelihood $p(\boldsymbol{x})$ is optimized using the ELBO (Kingma & Welling, 2013):
@@
$\log p(\boldsymbol{x})\geq\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x})}\left[\log p_{\theta}(\boldsymbol{x}\mid\boldsymbol{z})\right]-\operatorname{KL}(q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x})\|p(\boldsymbol{z})).$ (1)
@@
The ELBO balances two terms: the reconstruction term, which ensures the model accurately reconstructs the data from the latent representation, and the Kullback-Leibler (KL) divergence, which regularizes the learned latent distribution $q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x})$ to remain close to some pre-defined prior $p(\boldsymbol{z})$.
@@
Conditional VAEs (CVAEs) (Sohn et al., 2015) extend this framework to conditional distributions $p(\boldsymbol{x}\mid\boldsymbol{c})$, where $\boldsymbol{c}$ represents auxiliary information. In $\beta$-VAEs (Higgins et al., 2017), the KL divergence term is scaled by a hyperparameter $\beta$, leading to a modified ELBO as follows:
@@
$\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x},\boldsymbol{c})}\left[\log p_{\theta}(\boldsymbol{x}\mid\boldsymbol{z},\boldsymbol{c})\right]-\beta\operatorname{KL}(q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x},\boldsymbol{c})\|p(\boldsymbol{z})).$ (2)
@@
The choice of $\beta$ governs the trade-off between disentangling latent representations and reconstruction fidelity. While $\beta>1$ emphasizes disentanglement, in our setting, we use $\beta<1$ to prioritize accurate reconstruction over latent disentanglement.
To enhance flexibility, and reduce over-regularization we adopt the variational mixture of posteriors (VAMP) prior (Tomczak & Welling, 2018), which models the prior $p(\boldsymbol{z})$ through a learnable mixture of variational posteriors:
$p_{\phi,\omega}(\boldsymbol{z})=\frac{1}{K}\sum_{k=1}^{K}q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x}_{k},\boldsymbol{c}_{k}),$ (3)
where $\omega=\{\boldsymbol{x}_{k},\boldsymbol{c}_{k}\}^{K}_{k=1}$ are learned parameters that can be thought of as hyperparameters of the prior replacing actual inputs to the encoder.
@@ 2.2 Causal Models
In this paper, we employ the notation of structural causal models (SCM) for describing causal data generating processes. For a detailed definition, we refer the reader to (Bareinboim et al., 2022). Specifically, we treat a causal model $\mathcal{M}$ as a generative process over $d$ variables $\boldsymbol{V}=\{V_{1},...,V_{d}\}$. The model is causal in the sense that each variable is the direct cause of a subset of $\boldsymbol{V}$. Specifically, any variable $V_{j}\in\boldsymbol{V}$ is determined by its direct causes with $V_{j}\leftarrow f_{j}(Pa_{V_{j}},U_{j})$, where $f_{j}$ is an arbitrary function of the direct causes $Pa_{V_{j}}$ (representing parent variables of $V_{j}$) and a noise term $U_{j}$. All causal relations combined form the causal graph $G^{\mathcal{M}}$. Each model $\mathcal{M}$ induces a joint distribution $p_{\mathcal{M}}(\boldsymbol{V})$, called the observational distribution.
@@
In a causal model, performing a so-called intervention $do(V=v)$, manipulates $\mathcal{M}$ such that the target variable $V$ is forced to take on value $v$, regardless of $V$’s causes. Such an intervention results in an intervened model that we denote $\mathcal{M}_{do(V=v)}$ or $\mathcal{M}_{do(V)}$ when $v$ is clear from the context. Analogously, we denote the variables of $\mathcal{M}_{do(V=v)}$ as $\boldsymbol{V}_{do(V=v)}$. In this work, $v$ is one possible intervention value for which the training data contains samples. Similarly to its observational counterpart, $\mathcal{M}_{do(V)}$ induces a joint distribution $p_{\mathcal{M}}(\boldsymbol{V}\mid do(V))$ that we call the interventional distribution.
The task of distributional causal effect estimation is to identify $p_{\mathcal{M}}(\boldsymbol{V}\mid do(V))$ from an observational dataset $\boldsymbol{D}^{\mathcal{M}}=\boldsymbol{x}^{\mathcal{M}}_{1:N}$, where $N$ is the number of samples and $\boldsymbol{x}^{\mathcal{M}}_{n}\sim p_{\mathcal{M}}(\boldsymbol{V})$. In general, estimating this effect is not possible without additional assumptions or interventional data (Bareinboim et al., 2022). This work aims at alleviating this challenge by using interventional data at training time and exploiting the learned, amortized information at inference time without needing interventional data or causal graphs.

### v2

# ACTIVA: Amortized Causal Effect Estimation via Transformer-based Variational Autoencoder

@@ Abstract
Predicting the distribution of outcomes under hypothetical interventions is crucial across healthcare, economics, and policy-making. However, existing methods often require restrictive assumptions, and are typically limited by the lack of amortization across problem instances. We propose ACTIVA, a transformer-based conditional variational autoencoder (VAE) architecture for amortized causal inference, which estimates interventional distributions directly from observational data without. ACTIVA learns a latent representation conditioned on observational inputs and intervention queries, enabling zero-shot inference by amortizing causal knowledge from diverse training scenarios. We provide theoretical insights showing that ACTIVA predicts interventional distributions as mixtures over observationally equivalent causal models. Empirical evaluations on synthetic and semi-synthetic datasets confirm the effectiveness of our amortized approach and highlight promising directions for future real-world applications.
@@ 1 Introduction
Understanding the causal effects of an action by pure observation is central in many domains including healthcare (Shi and Norgeot 2022), economics (Panizza and Presbitero 2014), and finance (Kumar et al. 2023). Yet, estimating these effects is constraint by inherent limitations regarding the conclusions that can be drawn from observations only (Bareinboim et al. 2022). Specifically, looking only at observational data leads to ambiguities over which causal effect a certain action will have.
To address these ambiguities, current approaches to causal inference often rely on strong assumptions(see 7 for more detail). An alternative road to deal with these ambiguities is to make them explicit in the form of uncertainties. For example providing a interval or distribution of plausible causal effects has been shown to be feasible and useful in practice  (Maathuis et al. 2009; Balazadeh et al. 2025). Such a prediction allows for a clearer picture on what effects to expect and can lead to actionable inferences if e.g. the range of effects lies within an interval that leads to the same actions.
Although having a distribution over causal effects is a promising direction, it only represents uncertainty over this specific point-estimate. A natural extension of this is to look at the overall distribution shift of the causal factors after performing an action.Additionally to answering cause-effect questions, such a distribution can provide further task-relevant insights such as multimodality and hence prevent erroneous average effect predictions.
Recent work further emphasizes the reuse of learned knowledge of diverse causal tasks across different queries to improve inference efficiency by learning one model for estimation of different problem instances (Löwe et al. 2022; Lorch et al. 2022; Scetbon et al. 2024; Sauter et al. 2024; Mahajan et al. 2024; Annadani et al. 2025; Robertson et al. 2025; Balazadeh et al. 2025; Ma et al. 2025). Results from these approaches suggest that we can effectively amortize over datasets coming from different causal models for various downstream tasks.
@@
In this paper, we address above restrictions by proposing a causal encoder-decoder architecture called ACTIVA. During training, the encoder maps the interventional data to a latent space conditioned on observational data and an interventional query, and the decoder transforms the encoded causal information into interventional distribution. At inference time, we condition our prior on the observational data alone to obtain a distribution over plausible causal models which the decoder transforms into a mixture over these models. We provide a theoretical analysis on the specific nature of our latent and decoded space. In empirical evaluations, we show that our architecture can successfully recover causal information at inference time even on novel instances. This allows for the potential zero-shot transfer from simulated scenarios to the real-world ones without knowing the causal relations neither during, avoiding common pitfalls when relying on such graphs (Poinsot et al. 2024). Below we provide a detailed list of our main contributions:
@@
  • We provide a conditional variational autoencoder model (ACTIVA) designed for amortized causal distribution shift estimation. Our model takes an observational dataset and a query intervention as input, predicts a latent representation of the dataset and outputs an estimate of the respective interventional distribution.
  • We analyze the proposed model theoretically, leading to rare fundamental insights about amortized causal inference. The result emphasizes that amortized causal inference is inherently limited by identifiability from the data. Nevertheless it can recover shift distributions up to a mixture of shifts from observationally equivalent models.
@@
  • We implement an instantiation of ACTIVA and show empirically that it predicts interventional distributions given observational data and interventional queries both on synthetic and semi-synthetic data.
@@
Overall, our work highlights the significance of amortized causal inference as a tool to overcome traditional hurdles in distributional causal effect estimation. [^1] The code to our model is available at https://anonymous.4open.science//r/Amortized˙Interventional˙Distribution˙Estimation-E182/
2 Background and Notation
Next, we introduce the notation and the formal details for $\beta$-VAEs and causality.
2.1 Conditional $\beta$-VAEs
@@
Variational Autoencoders (VAEs) are generative models that approximate the joint distribution $p_{\theta}(\boldsymbol{x},\boldsymbol{z})=p_{\gamma}(\boldsymbol{x}|\boldsymbol{z})p_{\eta}(\boldsymbol{z})$, where $\boldsymbol{z}$ are latent variables governing the data $\boldsymbol{x}$ and $\theta=\{\gamma,\eta\}$ are the parameters determining the data generation. The marginal data likelihood $p_{\theta}(\boldsymbol{x})$ is optimized using the evidence lower bound (ELBO) (Kingma and Welling 2013):
@@
$\log p_{\theta}(\boldsymbol{x})\geq\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}|\boldsymbol{x})}\left[\log p_{\gamma}(\boldsymbol{x}|\boldsymbol{z})\right]-\operatorname{KL}(q_{\phi}(\boldsymbol{z}|\boldsymbol{x})\|p_{\eta}(\boldsymbol{z})).$ (1)
@@
Where $\phi$ parametrizes the encoding distribution $q$. The ELBO has a reconstruction term, which ensures the model accurately reconstructs the data from the latent representation, and a Kullback-Leibler divergence (KL) term, which regularizes the encoding distribution $q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x})$ to approximate some prior $p_{\eta}(\boldsymbol{z})$.
Conditional VAEs (CVAEs) (Sohn, Yan, and Lee 2015) extend this framework to allow for conditional generation with some auxiliary information $\boldsymbol{c}$. The generative process of CVAEs follows a conditional prior such that
$p_{\theta}(\boldsymbol{x},\boldsymbol{z}|\boldsymbol{c})=p_{\gamma}(\boldsymbol{x}|\boldsymbol{z},\boldsymbol{c})p_{\eta}(\boldsymbol{z}|\boldsymbol{c}).$ (2)
@@
In $\beta$-VAEs (Higgins et al. 2017), the KL term is scaled by a hyperparameter $\beta$, leading to a modified ELBO as follows:
@@
$\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}\mid\boldsymbol{x},\boldsymbol{c})}\left[\log p_{\gamma}(\boldsymbol{x}|\boldsymbol{z},\boldsymbol{c})\right]-\beta\operatorname{KL}(q_{\phi}(\boldsymbol{z}|\boldsymbol{x},\boldsymbol{c})\|p_{\eta}(\boldsymbol{z}|\boldsymbol{c})).$ (3)
@@
The choice of $\beta$ governs the trade-off between accurately reconstructing the data and disentangling the latent representations. While $\beta>1$ emphasizes disentanglement, in our setting, we use $\beta<1$ to prioritize accurate reconstruction.
@@ 2.2 Causal Models
In this paper, we employ the notation of structural causal models (SCM) for describing causal data generating processes. For a detailed definition, we refer the reader to (Bareinboim et al. 2022).
@@
We treat a causal model $\mathcal{M}$ as a generative process over $d$ variables $\boldsymbol{V}=\{V_{1},...,V_{d}\}$, denoting an assignment of these variables as $\boldsymbol{v}=\{v_{1},...,v_{d}\}$. The model is causal in the sense that each variable is the direct cause of a subset of $\boldsymbol{V}$. Specifically, any variable $V_{j}\in\boldsymbol{V}$ is determined by its direct causes with $V_{j}\leftarrow f_{j}(Pa_{V_{j}},U_{j})$, where $f_{j}$ is an arbitrary function of the direct causes $Pa_{V_{j}}$ (representing parent variables of $V_{j}$) and a noise term $U_{j}$. We call two models equivalent if all their parameters are equal. All causal relations combined form the causal graph $G^{\mathcal{M}}$. Each model $\mathcal{M}$ induces a joint distribution $p_{\mathcal{M}}(\boldsymbol{V})$, called the observational distribution. All models with the same observational distribution form a so-called Markov equivalence class (MEC).
@@
In a causal model, performing a so-called intervention $do(V=v)$, manipulates $\mathcal{M}$ such that the target variable $V$ is forced to take on value $v$, regardless of $V$’s causes. Such an intervention results in an intervened model that we denote $\mathcal{M}_{do(V=v)}$ or $\mathcal{M}_{do(V)}$ when $v$ is clear from the context. We denote the variables of $\mathcal{M}_{do(V=v)}$ as $\boldsymbol{V}\mid do(V=v)$. In this work, $v$ is one possible intervention value for which the training data contains samples. Similarly to its observational counterpart, $\mathcal{M}_{do(V)}$ induces a joint distribution $p_{\mathcal{M}}(\boldsymbol{V}\mid do(V))$ that we call the interventional distribution.
The aim of this work is to estimate the causally shifted distribution $p_{\mathcal{M}}(\boldsymbol{V}\mid do(V))$ from an observational dataset $\boldsymbol{D}^{\mathcal{M}}=\boldsymbol{v}^{\mathcal{M}}_{1:N}$, where $N$ is the number of i.i.d. samples and $\boldsymbol{v}^{\mathcal{M}}_{n}\sim p_{\mathcal{M}}(\boldsymbol{V})$. In general, identifying this distribution is not possible without additional assumptions or interventional data because of the ambiguities introduced by MECs (Bareinboim et al. 2022). This fundamental property of causality naturally extends to our work which we address in the theoretical portion of this work on data with identifiable effects.
3 ACTIVA
In this section, we outline a $\beta$-CVAE architecture for amortized causal distribution shift estimation called ACTIVA. In our setup, a dataset of observational samples and a query intervention serve as a conditional for the following generative process:
$p_{\theta}(\boldsymbol{V},\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))=p_{\gamma}(\boldsymbol{V}|\boldsymbol{z})p_{\eta}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))$ (4)
Intuitively, the data is modeled by a conditional prior that maps the observational data $\boldsymbol{D}^{\mathcal{M}}$ and query intervention $do(V)$ to a latent code $\boldsymbol{z}$. The latents then contain all information the model needs to reconstruct the data; in our case the samples from a post-interventional distribution $\boldsymbol{V}$.
To find parameters $\theta$, we use the conditional ELBO formulization for likelyhood maximization as in (3). Translated to our model, our overall learning objective becomes:
$\mathcal{L}(\theta,\phi)$ $=\displaystyle\mathop{\mathbb{E}}_{q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))}\left[\log p_{\gamma}(\boldsymbol{V}|\boldsymbol{z})\right]$
$-\operatorname{KL}(q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))\parallel p_{\eta}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))),$ (5)
where $q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))$ is the encoding distribution, $p_{\gamma}(\boldsymbol{V}|\boldsymbol{z})$ the decoder distribution, and $p_{\eta}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))$ the conditional prior.
To estimate the objective from data, we assume to have training data $\mathcal{\mathbf{D}}^{tr}=[\boldsymbol{v}^{\mathcal{M}_{do(V)}}_{n},\boldsymbol{D}^{\mathcal{M}},do(V)]_{j=0}^{J}$ with $J$ training examples pairing interventional samples with observational ones and the corresponding intervention. Where $\mathcal{\mathbf{D}}^{tr}_{j}\sim p_{tr}(\mathbf{D}^{tr})=\int p_{tr}(\mathbf{D}^{tr}\mid\mathcal{M})p_{tr}(\mathcal{M})d\mathcal{M}$ is an element from the training data distribution according to a pre-defined distributions of causal models $p_{tr}(\mathcal{M})$. This pre-defined distribution can be seen as the simulator that models causal models that we expect at inference time.
To enable gradient-based optimization, we use the reparameterization trick (Kingma and Welling 2013), allowing backpropagation through the stochastic sampling process of $q_{\phi}$. For our final objective, we aim at amortizing prediction over the training distribution of causal model, to successfully apply our model to data sets that were not in the training data. Optimizing our amortized objective then amounts to minimizing the expectation of our loss regarding the distribution of the training datasets.
$\min_{\theta,\phi}\quad\mathop{\mathbb{E}}_{\mathcal{\mathbf{D}}^{tr}_{j}\sim p_{tr}(\mathbf{D}^{tr})}\mathcal{L}(\theta,\phi)$ (6)
This objective ensures that our model can make inferences on datasets coming from the class of predefined causal models, even if the specific dataset was not in the training set. It furthermore implies that the more general the training distribution of data sets is, the more the model will generalize to new data sets during inference as has also been argued in (Montagna et al. 2024).
4 Theoretical Analysis
In this section, we show that inference with the model specified above gives us a mixture of all causal shift distributions that are consistent with the observational data. We start with two assumptions.
Assumption 1:
$q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))$ correctly identifies the causal model, i.e., $q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))=q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M^{\prime}}},do(V))\Rightarrow\mathcal{M}=\mathcal{M}^{\prime}$.
This assumption implies that the models in the training data $\mathbf{D}^{tr}$ must be identifiable from the observational and interventional data. Furthermore, we assume
Assumption 2:
The decoder is injective, i.e., for two inputs $\boldsymbol{z}\neq\boldsymbol{z^{\prime}}\Rightarrow dec_{\gamma}(\boldsymbol{z})\neq dec_{\gamma}(\boldsymbol{z^{\prime}})\Rightarrow p_{\gamma}(\boldsymbol{V}|\boldsymbol{z})\neq p_{\gamma}(\boldsymbol{V}|\boldsymbol{z^{\prime}})$.
Next, we show that the encoder distribution must place all probability mass on a single point.
Definition 1 (Dirac distribution, informal):
Let the Dirac distribution be defined via its density as $\delta_{z_{k}}(x)=\begin{cases}\infty\textit{ if }x=z_{k}\\ 0\textit{ otherwise }\end{cases}$ such that $\int_{-\infty}^{\infty}\delta_{z_{k}}(x)dx=1$. Such distribution places all probability on one point.
Lemma 1:
Under Assumptions 1-2, $q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))$ is a Dirac distribution.
Proof: Take any two equivalent models $\mathcal{M}=\mathcal{M^{\prime}}$ and two samples $z\sim q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V)),z^{\prime}\sim q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M^{\prime}}},do(V))$. Suppose that the two samples differ, i.e. $z\neq z^{\prime}$, then we know from Assumption 2 that $p_{\gamma}(\boldsymbol{V}|\boldsymbol{z})\neq p_{\gamma}(\boldsymbol{V}|\boldsymbol{z^{\prime}})$. But since we know that they represent the same model (Assumption 1), the two resulting distributions over $\boldsymbol{V}$ must be the same, which leads to a contradiction. Therefore, $z$ and $z^{\prime}$ must be the same. Accordingly, $q_{\phi}(\boldsymbol{z}|\boldsymbol{V},\boldsymbol{D}^{\mathcal{M}},do(V))$ must always yield the same sample, and the proof is complete.
Note that taking Lemma 1 and Assumption 1 together let’s us interpret $z_{k}$ as the latent representation of the causal model $\mathcal{M}^{k}$ and intervention query.
Next, we investigate what the learned prior will look like under the above assumptions. This will help us to further investigate the estimated interventional distribution.
Note that for the amortized loss in (6), we can rearrange the loss computation as follows by sorting the training samples accordingly: $\frac{1}{J}\sum_{\mathbf{D}^{tr}}\mathcal{L}=\frac{1}{J}\sum_{I}\sum_{O}\sum_{\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}}\mathcal{L}$, where $J$ is the number of datapoints, $I$ is the set of possible interventions, $O$ is a partition over the datasets according to their Markov equivalence class and $\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}$ is the set of training examples within the equivalence class with representative $\mathcal{M}^{o}$, s.t. $\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}=\{[\boldsymbol{v}^{\mathcal{M}_{do(V)}}_{n},\boldsymbol{D}^{\mathcal{M}}]_{j}\in\mathbf{D}^{tr}:p_{\mathcal{M}}(\boldsymbol{D}^{\mathcal{M}})=p_{\mathcal{M}^{o}}(\boldsymbol{D}^{\mathcal{M}^{o}})\}$. In other words, $\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}$ contains all data-points where the probability of generating such observational data-points by $\mathcal{M}$ is the same as the one by $\mathcal{M}^{o}$.
This decomposition is necessary to derive the alternative form of the ELBO according to (Hoffman and Johnson 2016) as described next. We define in the following $\bar{q}_{\phi}(\boldsymbol{z}|\mathbf{D}^{tr}_{[\mathcal{M}^{o}]},do(V)):=\frac{1}{J}\sum_{\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}}q_{\phi}(\boldsymbol{z}|\mathbf{D}^{tr}_{j})$ which is a uniform mixture of Dirac distributions as per Lemma 1 with $\pi_{c}=\frac{1}{J}$ for all component weights $\pi_{c}$ by construction. In other words, $\bar{q}_{\phi}$ is the distribution that averages over all encoder distributions that identify models that are observationally equivalent to $\boldsymbol{D}^{\mathcal{M}^{o}}$. Treating the sample index $j$ as random variable, we can further define $q(j,\boldsymbol{z}):=\frac{1}{J}q(\boldsymbol{z}|\mathbf{D}^{tr}_{j})$. Finally, this allows us to decompose the $\operatorname{KL}$ term in our loss into
$\frac{1}{J}\sum_{\mathbf{D}^{tr}_{[\mathcal{M}^{o}]}}\operatorname{KL}(q_{\phi}(\boldsymbol{z}|\mathcal{\mathbf{D}}^{tr}_{j})\parallel p_{\eta}(\boldsymbol{z}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{j}))$
$=\operatorname{KL}(\bar{q}_{\phi}(\boldsymbol{z}|\mathbf{D}^{tr}_{[\mathcal{M}^{o}]},do(V))||p_{\eta}(\boldsymbol{z}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{o}))$
$\quad+\mathbb{I}_{q(j,\boldsymbol{z})}[j,\boldsymbol{z}],$ (7)
where $\mathbb{I}_{q(j,\boldsymbol{z})}[j,\boldsymbol{z}]$ is the mutual information.
Since the resulting $\operatorname{KL}$ term is the only term that involves the prior, we focus our further analysis only on that term and disregard the mutual information. To ensure the $\operatorname{KL}$ is well-defined, we additionally make the following assumption.
Assumption 3:
The prior is a finite mixture of Dirac distributions which has zero mass, where $\bar{q}_{\phi}$ has zero mass, i.e., $p_{\eta}(\boldsymbol{z}|\boldsymbol{D}^{\mathcal{M}},do(V))=\sum_{C}\rho_{c}\delta_{\boldsymbol{z}_{c}}(z)$ s.t. $\forall\boldsymbol{z}_{c}:\bar{q}_{\phi}(\boldsymbol{\boldsymbol{z}_{c}}|\boldsymbol{D}^{[\mathcal{M}^{o}]},do(V))=0\Rightarrow p_{\eta}(\boldsymbol{z}_{c}|\boldsymbol{D}^{\mathcal{M}},do(V))=0$.
Intuitively, this implies that we have prior knowledge on which models do not fall into the observational equivalence class of the input dataset.
Proposition 1:
The prior $p_{\eta}(\boldsymbol{z}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{o})$ becomes equal to $\bar{q}_{\phi}(\boldsymbol{z}|\mathbf{D}^{tr}_{[\mathcal{M}^{o}]},do(V))\forall o\in O$ when the ELBO is minimized.
Proof: As argued above, the only term in our affecting $p_{\eta}$ is the $\operatorname{KL}$ term in (4) which needs to become zero to be minimized for each $o$. As per Assumption 3 and the definition of $\bar{q}_{\phi}$, this term becomes $\sum_{C}\pi_{c}\log\frac{\pi_{c}}{\rho_{c}}$ which is 0 exactly when $\forall c:\pi_{c}=\rho_{c}=\frac{1}{J}$. Hence, $p_{\eta}(\boldsymbol{z}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{o})=\bar{q}_{\phi}(\boldsymbol{z}|\mathbf{D}^{tr}_{[\mathcal{M}^{o}]},do(V))$ when (6) is minimized.
Lastly, we show that our method recovers a mixture of post-interventional distributions consistent with the observational data.
Proposition 2:
The conditional generative distribution $p_{\theta}(\boldsymbol{V}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{j})$ is a mixture of distributions $\sum_{\mathcal{M}^{k}\in[\mathcal{M}^{j}]}\rho_{k}p(\boldsymbol{V}|\boldsymbol{D}^{\mathcal{M}^{k}},[do(V)]_{j})$.
Proof: Recall the generative process (4). Specifically, we are interested in the marginal $p_{\theta}(\boldsymbol{V}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{j})$Because of Lemma 1 and Proposition 1 we can write the marginal as a sum over the latents $\boldsymbol{z}_{k}\in[\boldsymbol{z}_{j}]$ that represent models $\mathcal{M}^{k}\in[\mathcal{M}^{j}]$ yielding $\sum_{[\boldsymbol{z}_{j}]}p_{\gamma}(\boldsymbol{V}|\boldsymbol{z}_{k})p_{\eta}(\boldsymbol{z}_{k}|[\boldsymbol{D}^{\mathcal{M}},do(V)]_{j})$. Since $p_{\eta}$ is a uniform mixture, this is equivalent to $\sum_{[\boldsymbol{z}_{j}]}\rho_{k}p_{\gamma}(\boldsymbol{V}|\boldsymbol{z}_{k})=\sum_{\mathcal{M}^{k}\in[\mathcal{M}^{j}]}\rho_{k}p_{\theta}(\boldsymbol{V}|\boldsymbol{D}^{\mathcal{M}_{k}},[do(V)]_{j})$ since $\boldsymbol{z}_{k}$ encodes the model-query pair.
With this, we showed that ACTIVA learns to estimate the causal distribution shift from observational data and a query intervention, up to a mixture of observationally equivalent models. Our predicted distributions explicitly encode uncertainty about the data-generating causal model. Furthermore, these results are, to the best of our knowledge, alongside with results from Balazadeh et al. (2025) on uncertainty quantification in the identifiable data case, among the first steps to build a theory around amortized causal inference. Figure 1 shows the intuition behind our result.
4.1 Practical Considerations
Implementing our theoretical framework (Section 4) in practice involves several approximations. First, directly using Dirac distributions within a VAE framework is impractical due to undefined gradients. We thus approximate Dirac distributions with multivariate normal (MVN) distributions, using a fixed variance and zero covariance, which places each causal model in a distinct region of latent space. The prior is modeled as a Gaussian mixture model (GMM), naturally satisfying Assumption 3 and ensuring mathematical well-definedness of the KL divergence without explicit knowledge of latent locations.
Second, although our decoder (Section 5.4) violates injectivity (Assumption 2), empirical results suggest that this relaxation remains effective. Nonetheless, injective alternatives may offer a better theoretical alignment and represent a promising direction for future work.
Finally, since the prior is a GMM, obtaining closed-form solutions for interventional distributions at inference is intractable. We instead approximate the distribution by sampling 50 latent variables from the prior, decoding them, and constructing a uniform mixture. Despite these relaxations, our experiments demonstrate robust and practical estimation of interventional distributions.
Figure 1: Given observational data $\boldsymbol{D}^{\mathcal{M}}$ and a query intervention $do(X=x)$, the encoder maps to a latent point and the prior to a mixture of latent points. Samples $\boldsymbol{z}$ from this latent space are passed to the decoder, which estimates the interventional distribution $p_{\mathcal{M}}(\boldsymbol{V}\mid do(X=x))$ as a Gaussian mixture. Dashed lines indicate training-time information flow, solid lines represent inference-time.